Let z=5 where x, y, z E R. Prove that z? +z2+z?>
Answer #1

Let z=5 where x, y, z E R. Prove that z? +z2+z?>

Answers let n+y+z =5, X, Y, Z ER To show, x² + y² + Z² = ² Gбово, 144+2 = 5 >> KAY + 2 55 det k = 5/ -- x+y+z = K => * Tig +Addion these three relation we get - ܙ . . ܛܠܢ +: ܝ gi - z ; 3x3 [: ܐܪܨܪܐ ]_- 9 ** +z: 2 322 :*(%; [k - %) : 1 - z ܀ ܨܶܢ x